Circular No. 8365

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2004 A2, A3, B3–B10, C3–C8, D2–D4, E1–E4, F5 (SOHO)

Further to IAUC 8358, K. Battams reports measurements for additional near-sun comets found on SOHO coronagrapic images (and K. Mc-Gleam for C/2004 B4 and B5). Given below are comet designation, the date/time and position for the first available observation, the instrument(s) in which the comet was observed, the person who first found the comet on SOHO website images (JS = J. Sachs; TH = T. Hoffman; XZ = X.-m. Zhou; SH = S. Hönig; JD = J. Danaher; MB = M. Boschat; RK = R. Kracht; DC = D. Chen; KC = K. Černis), and the MPECs containing the reductions by R. Kisala and orbital elements by B. G. Marsden. All are Kreutz sungrazers — except C/2004 A3 and B3, which belong to the Kracht group, and C/2004 E2, which belongs to no known group. Note that the finder of C/2004 D3 was given erroneously on MPEC 2004-D14.

Comet 2004 UT α_{2000} δ_{2000} Inst. F MPEC
C/2004 A2 Jan. 9.821 19^{h}44.^{m}7 −27^{o}15′ C3/2 JS 2004-M04
C/2004 A3  15.996 19 54.3 −20 28 C2  TH 2004-M04
C/2004 B3  18.163 20 03.6 −19 57 C2  XZ 2004-M04
C/2004 B4  18.279 20 14.3 −23 21 C3  JS 2004-M04
C/2004 B5  19.446 20 19.7 −22 30 C3  SH 2004-M04
C/2004 B6  21.221 20 24.7 −21 54 C3  JD 2004-M10
C/2004 B7  21.363 20 29.3 −22 52 C3  MB 2004-M10
C/2004 B8  24.971 20 41.3 −20 59 C3  TH 2004-M10
C/2004 B9  25.738 20 48.7 −21 07 C3  MB 2004-M10
C/2004 B10  30.138 21 04.3 −19 22 C3  SH 2004-M10
C/2004 C3 Feb. 3.779 21 27.7 −19 05 C3  RK 2004-M13
C/2004 C4  8.071 21 38.7 −16 17 C3  XZ 2004-M13
C/2004 C5  8.929 21 40.4 −15 48 C3  TH 2004-M13
C/2004 C6  10.071 21 47.7 −15 44 C3  TH 2004-M13
C/2004 C7  13.571 22 04.3 −15 44 C3  RK 2004-M13
C/2004 C8  14.404 22 02.7 −14 05 C3  DC 2004-M13
C/2004 D2  25.013 22 54.6 −10 42 C3  TH 2004-M14
C/2004 D3  25.113 22 56.2 −10 42 C3  XZ 2004-M14
C/2004 D4  26.891 23 07.5 −10 19 C3/2 JS 2004-M14
C/2004 E1 Mar. 9.363 23 33.5 − 4 52 C3  KC 2004-M42
C/2004 E2  12.314 23 24.2 − 3 52 C2  RK 2004-M42
C/2004 E3  14.821 23 57.8 − 2 58 C3  RK 2004-M42
C/2004 E4  15.821 0 02.5 − 2 36 C3/2 SH 2004-M42
C/2004 F5  17.405 0 09.6 − 2 12 C3/2 RK 2004-M43
2004 July 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green